A boat goes 240 miles downriver in the same time it can go 160 miles upriver. The speed of the current is 5mph. Find the speed of the boat in still water.
Answer by ewatrrr(24785)

 
Hi
Note: D = rt Or t = D/r
A boat goes 240 miles downriver in the same time it can go 160 miles upriver. The speed of the current is 5mph. Find the speed 's' of the boat in still water.
240/(s+5mph) = 160/(s-5mph) Cross Multiplying to solve
240(s-5) = 160(s+5)
240s - 5*240 = 160s + 5*160
80s = 2000
s = 2000/80
s = 25mph, the speed of the boat in still water.
CHECKING our Answer***
240/30 = 160/20 = 8